If a train leaves at 8:00 and is traveling 60 miles an hour and a car leaves 4 hours later and is traveling at 80 miles an hour, when will the car catch up to the train
Answer by oberobic(2304)

Start with the basic distance equation:
d = r*t
.
Train:
d = 60*t
.
Car leaves 4 hours later going in the same direction:
d = 80*(t-4)
.
When the car overtakes the train, d = d.
.
60t = 80(t-4)
60t = 80t -320
-20t = -320
t = 16
.
In 16 hours the car will catch up.
8:00 a.m. + 16 hr = 24:00 = midnight
.
Check the distances to make sure they're the same.
60*16 = 960
The car travels for only 12 hours.
80*12 = 960
Correct.
.
Done.